you have just taken out a $15,000 car loan with a 7% APR, compounded monthly. The loan is for five years. When you make your first payment in one month, how much of the payment will go toward the principle of the loan and how much will go toward interest?
Answer #1

GIVEN:

You have just taken out a $15,000 car loan with a 7% APR, compounded monthly. The loan is for five years. When you make your first payment in one month.

TO FIND:

How much of the payment will go toward the principle of the loan and how much will go toward interest?

SOLUTION:

Payment=15000*(7%/12)/(1-1/(1+7%/12)^(12*5))

Interest:
=15000*7%/12=87.5

Principal:
=15000*(7%/12)/(1-1/(1+7%/12)^(12*5))-15000*7%/12=209.517978105242